# Fed-SB: A Silver Bullet for Extreme Communication Efficiency and Performance in (Private) Federated LoRA Fine-Tuning

### v1

@@ Abstract
Low-Rank Adaptation (LoRA) has become ubiquitous for efficiently fine-tuning foundation models. However, federated fine-tuning using LoRA is challenging due to suboptimal updates arising from traditional federated averaging of individual adapters. Existing solutions either incur prohibitively high communication cost that scales linearly with the number of clients or suffer from performance degradation due to limited expressivity. We introduce Federated Silver Bullet (Fed-SB), a novel approach for federated fine-tuning of LLMs using LoRA-SB, a recently proposed low-rank adaptation method. LoRA-SB optimally aligns the optimization trajectory with the ideal low-rank full fine-tuning projection by learning a small square matrix ($R$) between adapters $B$ and $A$, keeping other components fixed. Direct averaging of $R$ guarantees exact updates, substantially reducing communication cost, which remains independent of the number of clients, and enables scalability. Fed-SB achieves state-of-the-art performance across commonsense reasoning, arithmetic reasoning, and language inference tasks while reducing communication costs by up to 230x. In private settings, Fed-SB further improves performance by (1) reducing trainable parameters, thereby lowering the noise required for differential privacy and (2) avoiding noise amplification introduced by other methods. Overall, Fed-SB establishes a new Pareto frontier in the tradeoff between communication and performance, offering an efficient and scalable solution for both private and non-private federated fine-tuning. Our code is publicly available at https://github.com/CERT-Lab/fed-sb.
Fed-SB: A Silver Bullet for Extreme Communication Efficiency and Performance in (Private) Federated LoRA Fine-Tuning
Raghav Singhal*1, Kaustubh Ponkshe*1, Rohit Vartak2, Lav R. Varshney3, Praneeth Vepakomma1,4 1Mohamed bin Zayed University of Artificial Intelligence, 2Duke University, 3University of Illinois Urbana-Champaign, 4Massachusetts Institute of Technology *Equal contribution. Order decided by coin toss.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable generalization across a wide range of tasks (Achiam et al., 2023; Touvron et al., 2023; Team et al., 2023; Raffel et al., 2020). Fine-tuning (FT) remains the most effective approach for aligning LLMs to specific data distributions and reinforcing desired properties. However, as model sizes scale, full FT becomes increasingly prohibitive due to its substantial computational cost. To address this, parameter-efficient fine-tuning (PEFT) techniques, such as low-rank adaptation (LoRA, Hu et al. (2021)), have emerged as viable alternatives, offering a favorable trade-off between computational efficiency and performance. Variants of LoRA, including QLoRA (Dettmers et al., 2024), DoRA (Liu et al., 2024), AdaLoRA (Zhang et al., 2023), and LoRA-SB (Ponkshe et al., 2024), further refine this paradigm by optimizing memory efficiency, training dynamics, and generalization.
Federated learning (FL) is a popular method for training models in settings where data is siloed across multiple entities (Konečný et al., 2017; Kairouz et al., 2021; Bonawitz et al., 2019). Federated FT extends this paradigm by enabling large models, pre-trained on public data, to be efficiently adapted to private, distributed datasets without requiring clients to share their local data. Existing methods predominantly rely on LoRA-based techniques to learn client-specific adaptations (Zhang et al., 2024b). However, optimizing federated aggregation often involves tradeoffs between model performance (Sun et al., 2024) and communication efficiency (Wang et al., 2024; Singhal et al., 2025), necessitating careful design choices to balance these competing objectives.
@@
LoRA-SB (Ponkshe et al., 2024), a state-of-the-art approach, optimally simulates full fine-tuning in low-rank spaces by learning an $r\times r$ matrix between the low-rank adapters $\mathbf{A}$ and $\mathbf{B}$ while keeping other components fixed. This design reduces trainable parameters and enables better updates through its initialization strategy. Moreover, LoRA-SB demonstrates that this optimal approximation is not achievable with standard LoRA-based methods. LoRA-SB learns higher-rank updates with 2–4x greater rank than LoRA while requiring 45-90x fewer parameters. We propose Fed-SB, a federated variant of LoRA-SB, providing an ideal framework for (private) federated FT. Fed-SB overcomes limitations in LoRA-based federated FT while being significantly more computation- and communication-efficient. Notably, it enables exact and optimal aggregation by simply averaging the learnable matrix $\mathbf{R}$.
Differential privacy (DP) is a well-established framework for ensuring strong privacy guarantees (Dwork, 2006; Dwork et al., 2014), which is particularly crucial in federated settings. DP-SGD is a widely used privacy-preserving optimization method (Abadi et al., 2016), but its challenges are exacerbated in federated FT, where noise injected for privacy amplifies divergence across client models (Sun et al., 2024). Learning in DP-SGD is more effective when the number of learnable parameters is reduced, as the magnitude of noise added for privacy guarantees scales with the parameter count. Fed-SB mitigates this issue to yield improved performance, since it inherently has fewer learnable parameters and thus less noise injection. Furthermore, we show that Fed-SB avoids noise amplification introduced by other methods, further enhancing privacy-preserving learning.
Fed-SB pushes the performance vs communication cost Pareto frontier, offering an extremely efficient and scalable solution for both private and non-private federated FT, as shown in Figure 1. It consistently has superior performance while substantially reducing communication overhead than other methods. Our key contributions are summarized as follows.
@@
  • Fed-SB significantly reduces communication cost—by up to 230x compared to state-of-the-art methods—by requiring only an $r\times r$ matrix to be transmitted per aggregation.
  • We demonstrate that Fed-SB is particularly well-suited for privacy-preserving federated fine-tuning, as it minimizes noise by reducing the number of learnable parameters and leveraging linearity in the aggregate update.
  • Extensive experiments on four models across three diverse benchmarks show that Fed-SB consistently outperforms existing methods while drastically reducing communication overhead in both private and non-private federated settings, establishing a new Pareto frontier in federated fine-tuning.
@@
Learnable params. / client $\mathcal{O}((m+n)r)$ $\mathcal{O}((m+n)r)$ $\mathcal{O}((m+n)r)$ $\mathcal{O}(mr)$ $\mathcal{O}(r^{2})$
@@
No noise amplification
Privacy benefit (less params.)
@@
Table 1: Advantages of Fed-SB over various state-of-the-art federated fine-tuning methods ($c$ clients).
@@ 2 Preliminaries and Motivation
Federated Fine-Tuning. Given a pretrained weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, the objective in FT is to learn an update $\Delta\mathbf{W}$ for a given dataset. LoRA (Hu et al., 2021) remains the preferred method, where low-rank adapter matrices $\mathbf{A}\in\mathbb{R}^{r\times n}$ and $\mathbf{B}\in\mathbb{R}^{m\times r}$ are learned such that $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$. In federated learning, the dataset is distributed across $c$ clients, and the goal is to learn $\Delta\mathbf{W}$ without sharing local data with a central server. To achieve this, each client learns its own adapter matrices $\mathbf{A}_{i}$ and $\mathbf{B}_{i}$. The server aggregates these updates to refine $\mathbf{W}$, along with globally beneficial representations of $\mathbf{A}$ and $\mathbf{B}$, ultimately producing a shared aggregate model $\mathbf{W}^{\text{agg}}$. Next, each client continues the local FT process, followed by aggregation at the end of each round. This cycle repeats over multiple rounds. We summarize some of the state-of-the-art federated FT methods below.
Fed-IT (Zhang et al., 2024b) updates the adapters $\mathbf{A}$ and $\mathbf{B}$ using the standard FedAvg (McMahan et al., 2017) algorithm:
@@
FedEx-LoRA (Singhal et al., 2025) follows the same aggregation but introduces an additional error correction matrix $\mathbf{W}_{\text{err}}$ of rank $\min(cr,m,n)$:
@@
FLoRA (Wang et al., 2024) follows the same principle as FedEx-LoRA but achieves it by stacking the adapter matrices, and reinitializes them randomly at the end of each communication round.
FFA-LoRA (Sun et al., 2024) keeps $\mathbf{A}$ fixed while training (and aggregating) only $\mathbf{B}$ matrices.
@@
(Approximate) Differential Privacy. DP, introduced by Dwork (2006), is a widely adopted mathematical framework for privacy preservation. A randomized mechanism $\mathcal{M}:\mathcal{D}\rightarrow\mathcal{R}$, mapping a domain $\mathcal{D}$ to a range $\mathcal{R}$, satisfies $(\epsilon,\delta)$-differential privacy if, for any two adjacent inputs $d,d^{\prime}\in\mathcal{D}$ and any subset of outputs $S\subseteq\mathcal{R}$, the following holds:
@@
DP-SGD. DP-SGD (Abadi et al., 2016) is a privacy-preserving variant of stochastic gradient descent (SGD) designed to ensure DP during model training. It enforces privacy by clipping per-sample gradients to a fixed norm $C$ to limit their sensitivity and then adding isotropic Gaussian noise $\mathcal{N}\left(0,\sigma^{2}C^{2}\mathbf{I}\right)$, where $\sigma$ controls the noise magnitude. The cumulative privacy loss over iterations is quantified using the moments accountant (Wang et al., 2019) and Rényi DP (Mironov, 2017), which offer a tight bound on the final privacy parameter $\epsilon$.
@@
Standard federated averaging of individual LoRA adapters (FedIT, Zhang et al. (2024b)) introduces inexactness in aggregation, as the ideal update should be the average of client updates.
$\underbrace{\mathbf{W}_{0}+\frac{1}{c}\sum_{i=1}^{c}\mathbf{B}_{i}\times\frac{1}{c}\sum_{i=1}^{c}\mathbf{A}_{i}}_{\text{Vanilla aggregation in LoRA (FedIT)}}\\ \neq\underbrace{\mathbf{W}_{0}+\frac{1}{c}\sum_{i=1}^{c}(\mathbf{B}_{i}\mathbf{A}_{i})}_{\text{Ideal aggregation}}.$ (5)
@@
FFA-LoRA addresses inexact aggregation by keeping only $\mathbf{B}$ trainable while fixing $\mathbf{A}$ uniformly across clients. However, this comes at the cost of reduced expressivity and limits the benefits of jointly optimizing $\mathbf{A}$ and $\mathbf{B}$. As a result, performance degrades, as demonstrated previously (Singhal et al., 2025). This stems from two factors: suboptimal individual updates and the need for higher-rank adaptations. Freezing $\mathbf{A}$ leads to suboptimal updates, even in centralized training, where FFA-LoRA underperforms compared to LoRA. Additionally, Mahla et al. (2025) show that models trained using FFA-LoRA progressively deviate from the optimal hypothesis. Empirical evidence shows that the advantages of exactness are outweighed by the degradation caused by these factors.
Private Fine-Tuning. Pre-training on public data followed by FT on user-specific private data [^1] Although pre-training data may be publicly available, it often contains sensitive or proprietary information, raising privacy concerns. However, any privacy loss from pre-training has already occurred upon the model’s release. is a common approach for adapting models under privacy constraints (Yu et al., 2021; Tang et al., 2024). This two-stage process enhances performance in private learning while preserving user data privacy. FL naturally improves privacy by keeping data decentralized. However, even without direct data sharing, client-specific model updates can still leak sensitive information (Truong et al., 2021). Thus, developing privacy-preserving FT methods for FL is essential to ensure strong privacy guarantees while maintaining performance.
@@
$\Delta\mathbf{W}_{DP}-\Delta\mathbf{W}=\left(\mathbf{B}+\boldsymbol{\xi}_{B}\right)\left(\mathbf{A}+\boldsymbol{\xi}_{A}\right)-\mathbf{B}\mathbf{A}\\ =\boldsymbol{\xi}_{B}\mathbf{A}+\mathbf{B}\boldsymbol{\xi}_{A}+\boldsymbol{\xi}_{B}\boldsymbol{\xi}_{A}.$ (6)
The first-order noise term, $\boldsymbol{\xi}_{B}\mathbf{A}+\mathbf{B}\boldsymbol{\xi}_{A}$, is expected and occurs even in full FT with DP-SGD. However, the second-order noise term, $\boldsymbol{\xi}_{B}\boldsymbol{\xi}_{A}$, causes noise amplification, leading to further performance degradation in LoRA-based methods (Sun et al., 2024). This issue is exacerbated in FL, as individual client updates deviate even further from the ideal global update. FFA-LoRA avoids this problem by freezing $\mathbf{A}$, preventing the introduction of additional noise terms.
A Silver Bullet Indeed. The bilinear parameterization in LoRA introduces two key challenges: inexact aggregation and noise amplification. FedEx-LoRA/FLoRA addresses the inexactness issue by enabling exact aggregation, but at the cost of a communication overhead that scales prohibitively with the number of clients. FFA-LoRA mitigates inexact aggregation and excessive communication but sacrifices performance, as it operates in a low-rank space and has reduced expressivity. An ideal method would efficiently learn higher-rank updates while inherently enabling exact aggregation without increasing communication costs. However, any LoRA-based formulation that attempts to resolve all these challenges must inevitably trade off expressivity, ultimately compromising performance. We prove that LoRA-SB provides an optimal reparameterization of the updates, effectively overcoming all limitations of LoRA in both non-private and privacy-preserving federated settings.

### v2

@@ Abstract
Low-Rank Adaptation (LoRA) has become ubiquitous for efficiently fine-tuning foundation models. However, federated fine-tuning using LoRA is challenging due to suboptimal updates arising from traditional federated averaging of individual adapters. Existing solutions either incur prohibitively high communication cost that scales linearly with the number of clients or suffer from performance degradation due to limited expressivity. We introduce Federated Silver Bullet (Fed-SB), a novel approach for federated fine-tuning of LLMs using LoRA-SB, a recently proposed low-rank adaptation method. LoRA-SB optimally aligns the optimization trajectory with the ideal low-rank full fine-tuning projection by learning a small square matrix ($R$) between adapters $B$ and $A$, keeping other components fixed. Direct averaging of $R$ guarantees exact updates, substantially reducing communication cost, which remains independent of the number of clients, and enables scalability. Fed-SB achieves state-of-the-art performance across commonsense reasoning, arithmetic reasoning, and language inference tasks while reducing communication costs by up to 230x. In private settings, Fed-SB further improves performance by (1) reducing trainable parameters, thereby lowering the noise required for differential privacy and (2) avoiding noise amplification introduced by other methods. Overall, Fed-SB offers a state-of-the-art, efficient, and scalable solution for both private and non-private federated fine-tuning. Our code is publicly available at: https://github.com/CERT-Lab/fed-sb.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable generalization across a wide range of tasks (2, 49, 46, 40). Fine-tuning (FT) remains the most effective approach for aligning LLMs to specific data distributions and reinforcing desired properties. However, as model sizes scale, full FT becomes increasingly prohibitive due to its substantial computational cost. To address this, parameter-efficient fine-tuning (PEFT) techniques, such as low-rank adaptation (LoRA, (21)), have emerged as viable alternatives, offering a favorable trade-off between computational efficiency and performance. Variants of LoRA, including QLoRA (14), DoRA (32), AdaLoRA (60), and LoRA-SB (39), further refine this paradigm by optimizing memory efficiency, training dynamics, and generalization.
Federated learning (FL) is a popular method for training models in settings where data is siloed across multiple entities (26, 24, 7). Federated FT extends this paradigm by enabling large models, pre-trained on public data, to be efficiently adapted to private, distributed datasets without requiring clients to share their local data. Existing methods predominantly rely on LoRA-based techniques to learn client-specific adaptations (58). However, optimizing federated aggregation often involves tradeoffs between model performance (44) and communication efficiency (52, 43), necessitating careful design choices to balance these competing objectives.
@@
LoRA-SB (39), a state-of-the-art approach, optimally simulates full fine-tuning in low-rank spaces by learning an $r\times r$ matrix between the low-rank adapters $\mathbf{A}$ and $\mathbf{B}$ while keeping other components fixed. This design reduces trainable parameters and enables better updates through its initialization strategy. Moreover, LoRA-SB demonstrates that this optimal approximation is not achievable with standard LoRA-based methods. LoRA-SB learns higher-rank updates with 2–4x greater rank than LoRA while requiring 45-90x fewer parameters. We propose Fed-SB, a federated variant of LoRA-SB, providing an ideal framework for (private) federated FT. Fed-SB overcomes limitations in LoRA-based federated FT while being significantly more computation- and communication-efficient. Notably, it enables exact and optimal aggregation by simply averaging the learnable matrix $\mathbf{R}$.
Differential privacy (DP) is a well-established framework for ensuring strong privacy guarantees (17, 18), which is particularly crucial in federated settings. DP-SGD is a widely used privacy-preserving optimization method (1), but its challenges are exacerbated in federated FT, where noise injected for privacy amplifies divergence across client models (44). Learning in DP-SGD is more effective when the number of learnable parameters is reduced, as the magnitude of noise added for privacy guarantees scales with the parameter count. Fed-SB mitigates this issue to yield improved performance, since it inherently has fewer learnable parameters and thus less noise injection. Furthermore, we show that Fed-SB avoids noise amplification introduced by other methods, further enhancing privacy-preserving learning.
Fed-SB pushes the performance vs communication cost Pareto frontier, offering an extremely efficient and scalable solution for both private and non-private federated FT, as shown in Figure 1. It consistently has superior performance while substantially reducing communication overhead than other methods. Our key contributions are summarized as follows:
@@
  • Fed-SB consistently achieves state-of-the-art results while significantly reducing communication cost, by up to 230x, by requiring only an $r\times r$ matrix to be transmitted per aggregation.
  • We demonstrate that Fed-SB is particularly well-suited for privacy-preserving (federated) fine-tuning, as it minimizes noise by reducing the number of learnable parameters and leveraging linearity in the aggregate update.
  • Extensive experiments on $4$ models across $3$ diverse benchmarks show that Fed-SB consistently outperforms existing methods while drastically reducing communication overhead in both private and non-private federated settings, establishing a new Pareto frontier in federated fine-tuning.
Table 1: Advantages of Fed-SB over various SOTA federated fine-tuning methods ($c$ clients). Fed-SB achieves exact aggregation and high expressivity with extremely low communication cost - constant with the number of clients. In private settings, Fed-SB offers additional advantages by minimizing noise through reducing learnable parameters and leveraging linearity to avoid noise amplification.
@@
Learnable params. $\mathcal{O}((m+n)r)$ $\mathcal{O}((m+n)r)$ $\mathcal{O}((m+n)r)$ $\mathcal{O}(mr)$ $\mathcal{O}(r^{2})$
@@
No noise ampl.
Privacy (less params.)
@@ 2 Preliminaries and Motivation
Federated Fine-Tuning. Given a pretrained weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, the objective in FT is to learn an update $\Delta\mathbf{W}$ for a given dataset. LoRA (21) remains the preferred method, where low-rank adapter matrices $\mathbf{A}\in\mathbb{R}^{r\times n}$ and $\mathbf{B}\in\mathbb{R}^{m\times r}$ are learned such that $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$. In federated learning, the dataset is distributed across $c$ clients, and the goal is to learn $\Delta\mathbf{W}$ without sharing local data with a central server. To achieve this, each client learns its own adapter matrices $\mathbf{A}_{i}$ and $\mathbf{B}_{i}$. The server aggregates these updates to refine $\mathbf{W}$, along with globally beneficial representations of $\mathbf{A}$ and $\mathbf{B}$, ultimately producing a shared aggregate model $\mathbf{W}^{\text{agg}}$. Next, each client continues the local FT process, followed by aggregation at the end of each round. This cycle repeats over multiple rounds. We summarize some of the state-of-the-art federated FT methods below.
Fed-IT (58) updates the adapters $\mathbf{A}$ and $\mathbf{B}$ using the standard FedAvg (35) algorithm:
@@
FedEx-LoRA (43) follows the same aggregation but introduces an additional error correction matrix $\mathbf{W}_{\text{err}}$ of rank $\min(cr,m,n)$:
@@
FLoRA (52) follows the same principle as FedEx-LoRA but achieves it by stacking the adapter matrices, and reinitializes them randomly at the end of each communication round. FFA-LoRA (44) keeps $\mathbf{A}$ fixed while training (and aggregating) only $\mathbf{B}$ matrices.
@@
[…]
(Approximate) Differential Privacy. DP, introduced by (17), is a widely adopted mathematical framework for privacy preservation. A randomized mechanism $\mathcal{M}:\mathcal{D}\rightarrow\mathcal{R}$, mapping a domain $\mathcal{D}$ to a range $\mathcal{R}$, satisfies $(\epsilon,\delta)$-differential privacy if, for any two adjacent inputs $d,d^{\prime}\in\mathcal{D}$ and any subset of outputs $S\subseteq\mathcal{R}$, the following holds:
@@
$\mathbf{B}^{j+1}_{i}\leftarrow\frac{1}{k}\sum_{i=1}^{k}\mathbf{B}^{j}_{i},\mathbf{A}^{j+1}_{i}\leftarrow\frac{1}{k}\sum_{i=1}^{k}\mathbf{A}^{j}_{i},\mathbf{W_{0}}^{j+1}\leftarrow\mathbf{W_{0}}^{j}+\underbrace{\frac{1}{k}\sum_{i=1}^{k}(\mathbf{B}^{j}_{i}\mathbf{A}^{j}_{i})-\frac{1}{k}\sum_{i=1}^{k}\mathbf{B}^{j}_{i}\times\frac{1}{k}\sum_{i=1}^{k}\mathbf{A}^{j}_{i}}_{\text{Residual}}$ (6)
DP-SGD. DP-SGD (1) is a privacy-preserving variant of stochastic gradient descent (SGD) designed to ensure DP during training. It enforces privacy by clipping per-sample gradients to a fixed norm $C$ to limit their sensitivity and then adding isotropic Gaussian noise $\mathcal{N}\left(0,\sigma^{2}C^{2}\mathbf{I}\right)$, where $\sigma$ controls the noise magnitude. The cumulative privacy loss over iterations is quantified using the moments accountant (51) and Rényi DP (38), which offer a tight bound on the final privacy parameter $\epsilon$.
@@
Standard federated averaging of individual LoRA adapters (FedIT (58)) introduces inexactness in aggregation, as the ideal update should be the average of client updates.
$\underbrace{\mathbf{W}_{0}+\frac{1}{c}\sum_{i=1}^{c}\mathbf{B}_{i}\times\frac{1}{c}\sum_{i=1}^{c}\mathbf{A}_{i}}_{\text{Vanilla aggregation in LoRA (FedIT)}}\neq\underbrace{\mathbf{W}_{0}+\frac{1}{c}\sum_{i=1}^{c}(\mathbf{B}_{i}\mathbf{A}_{i})}_{\text{Ideal aggregation}}.$ (7)
@@
FFA-LoRA addresses inexact aggregation by keeping only $\mathbf{B}$ trainable while fixing $\mathbf{A}$ uniformly across clients. However, this comes at the cost of reduced expressivity and limits the benefits of jointly optimizing $\mathbf{A}$ and $\mathbf{B}$. As a result, performance degrades, as demonstrated previously (43). This stems from two factors: suboptimal individual updates and the need for higher-rank adaptations. Freezing $\mathbf{A}$ leads to suboptimal updates, even in centralized training, where FFA-LoRA underperforms compared to LoRA. Additionally, recent work (34) shows that models trained using FFA-LoRA progressively deviate from the optimal hypothesis. Empirical evidence shows that the advantages of exactness are outweighed by the degradation caused by these factors.
Private Fine-Tuning. Pre-training on public data followed by FT on user-specific private data [^1] Although pre-training data may be public, it often contains sensitive or proprietary information, raising privacy concerns. However, any privacy loss from pre-training has already occurred upon the model’s release. is a common approach for adapting models under privacy constraints (54, 45). This two-stage process enhances performance in private learning while preserving user data privacy. FL naturally improves privacy by keeping data decentralized. However, even without direct data sharing, client-specific model updates can still leak sensitive information (50). Thus, developing privacy-preserving FT methods for FL is essential to ensure strong privacy guarantees while maintaining performance.
@@
$\Delta\mathbf{W}_{DP}-\Delta\mathbf{W}=\left(\mathbf{B}+\boldsymbol{\xi}_{B}\right)\left(\mathbf{A}+\boldsymbol{\xi}_{A}\right)-\mathbf{B}\mathbf{A}=\boldsymbol{\xi}_{B}\mathbf{A}+\mathbf{B}\boldsymbol{\xi}_{A}+\boldsymbol{\xi}_{B}\boldsymbol{\xi}_{A}.$ (8)
The first-order noise term, $\boldsymbol{\xi}_{B}\mathbf{A}+\mathbf{B}\boldsymbol{\xi}_{A}$, is expected and occurs even in full FT with DP-SGD. However, the second-order noise term, $\boldsymbol{\xi}_{B}\boldsymbol{\xi}_{A}$, causes noise amplification, leading to further performance degradation in LoRA-based methods (44). This issue is exacerbated in FL, as individual client updates deviate even further from the ideal global update. FFA-LoRA avoids this problem by freezing $\mathbf{A}$, preventing the introduction of additional noise terms.
A Silver Bullet Indeed. The bilinear parameterization in LoRA introduces two key challenges: inexact aggregation and noise amplification. FedEx-LoRA/FLoRA addresses the inexactness issue by enabling exact aggregation, but at the cost of communication overhead that scales prohibitively with the number of clients. FFA-LoRA mitigates inexact aggregation and excessive communication but sacrifices performance, as it operates in a low-rank space and has reduced expressivity. An ideal method would efficiently learn higher-rank updates while inherently enabling exact aggregation without increasing communication costs. However, any LoRA-based formulation that attempts to resolve these challenges must inevitably trade off expressivity, ultimately compromising performance. We prove that LoRA-SB provides an optimal reparameterization of the updates, effectively overcoming all limitations of LoRA in both non-private and privacy-preserving federated settings.